OpenStudy (anonymous):

If f(x) = log3 (x + 1), what is f^−1 (2)? 1 8 10 27

OpenStudy (freckles):

f(x)=y if f has an inverse then f^{-1}(y)=x so if you have f^{-1}(2)=x then f(x)=2 so solve f(x)=2 for x that is solve log_3(x+1)=2

OpenStudy (anonymous):

I still dont understand

OpenStudy (anonymous):

Plug in 2 for x? So log_3 ( 2+1 ) = 2?

OpenStudy (freckles):

no

OpenStudy (freckles):

solve for x

OpenStudy (freckles):

\[\log_3(x+1)=2\]

OpenStudy (freckles):

recall \[3^{\log_3(a)}=a \text{ assuming } a>0\]

OpenStudy (freckles):

so that is you have \[3^{\log_3(x+1)}=3^2 \]

OpenStudy (freckles):

using the thought I just mentioned you have x+1=3^2 solve for x

OpenStudy (anonymous):

So the answer would be 8?

OpenStudy (freckles):

yes 9-1 is 8